# Interlayer exciton condensates between second Landau level orbitals in double bilayer graphene

### v1

@@ Abstract
Electrons moving in two-dimensional (2D) space and under a strong magnetic field fall into discrete Landau levels (LLs), indexed by an orbital quantum number $N=0,1,2$..., with the corresponding spatial wavefunction becoming increasingly nodal as $N$ increases. The ultraflat electronic dispersion creates an ideal environment for Coulomb interactions to dominate and induce various many-body ground states, such as the Wigner crystal [1], charge density wave [2], and most notably, the fractional quantum Hall (FQH) effect [3], where quasiparticles with fractional charge and statistics emerge. The orbital character is critical to interaction physics because the Haldane pseudopotential for LL wavefunctions varies with $N$ [4]. In particular, even-denominator FQH states at half-filling have been predominantly observed in the $N=1$ LL [5, 6, 7, 8, 9, 10], and are believed to be described by a class of non-Abelian states including the Moore-Read Pfaffian state [11]. There is also numerical evidence showing that composite fermion pairing is favored here due to the softened Coulomb repulsion, unlike in $N=0$ [12]. Experimentally, access to different LLs effectively serves as a powerful tuning knob for electron-electron interactions, fostering exotic many-body phases.
@@
Quantum Hall (QH) physics can be further enriched by placing two electrically isolated, parallel layers of 2D electron systems close to each other so that electrons in one layer interact solely through Coulomb interaction with those in the other. This double-layer structure enables more complex multi-component states by introducing a new pseudospin — the layer index, and provides unique experimental access to coherent QH states by allowing for probing the two spatially separated pseudospin components, something impossible for a spin or valley based multi-component state. In particular, the multi-component Halperin (111) state  [13] composed of layer pseudospins maps to an interlayer exciton condensate (EC) formed from electrons and holes in the partially filled LLs in the opposite layers [14, 15]. Such interlayer ECs were first observed in GaAs double quantum wells [16, 17], and later realized in double layer structures of monolayer or Bernal-stacked bilayer graphene (BLG) [18, 19, 20, 21]. The exceptional flexibility and tunability of these double-layer graphene structures further enabled observation of the BEC-BCS crossover [22], novel interlayer FQH states [20, 21] and exciton solids [23].
@@
However, all of these interlayer states have only been observed when both layers are in the $N=0$ LL; no such states have been detected between the $N=1$ LLs in graphene double-layer structures. Since $N=1$ can host a different hierarchy of single-layer FQH states, it could likewise support qualitatively new interlayer states in double-layer systems. Furthermore, once an EC between two $N=1$ LLs is realized, the transition between this interlayer coherent EC and individual-layer even-denominator FQH states could be a highly nontrivial topological phase transition [24, 25]. Previously, there have been reports of gapped states in 2H-stacked bilayer $\mathrm{WSe_{2}}$ [26] or twisted double bilayer graphene [27] that are consistent with interlayer ECs in higher LLs. However, while the absence of a spacer helps achieve strong interlayer coupling, it prevents probing each layer pseudospin separately, making it impossible to conclusively identify the interlayer coherence of EC states. In this Letter, we present the first observation of quantized Coulomb-drag signals between the $N=1$ LLs in a double bilayer graphene device, providing key signatures of EC formation in higher LLs.
@@
Fig. 1a shows the image of the device, and Fig. 1b illustrates its layer configuration. The device fabrication processes are similar to those in our previous studies [20, 21]. Briefly, two BLG layers are stacked on top of each other separated by a 2.5 nm thick hBN spacer. They are further encapsulated by layers of hBN and graphite from both top and bottom, with the graphite layers acting as local top and bottom gates. Separate gold contacts are made to the top and bottom BLGs where they do not overlap. To ensure good equilibrium between the gold contacts and the channel region, we utilize the global silicon back gate and fabricate additional aluminum oxide dielectric layers with patterned gold contact gates on top of the heterostructure to heavily dope the top and bottom graphene leads.
@@
The phase diagram of BLG at high magnetic fields is determined by the displacement field and carrier density. When higher-order energy anisotropies are ignored, the zero-energy LL of BLG consists of eight quasi-degenerate flavors: spin up and down, valley $K$ and $K^{\prime}$, and an orbital quantum number, $N=0$ and 1 [28]. Fig. 1c illustrates the latter two flavors. The $N=0$ orbital in BLG has a LL wavefunction similar to the lowest LL in a conventional quadratic-band system like GaAs. By contrast, the $N=1$ orbital in BLG is an admixture of the conventional lowest and second LL wavefunctions [28]. Depending on the valley flavor, these wavefunctions are layer-polarized within BLG, as shown in Fig. 1c. Previous studies on QH states in BLG have shown that next-order energy anisotropies and Coulomb interaction induce an energy hierarchy among the eight flavors, leading to their sequential filling as carrier density increases [29, 9, 8]. A displacement field further modifies the hierarchy by shifting the relative potential of the two graphene layers within the BLG and alters the energetics of the eight flavors due to their varying degrees of layer polarization.
@@
In our experiment, by applying DC voltages to the top and bottom graphite gates, $V_{\text{TG}}$ and $V_{\text{BG}}$, as well as between the two BLG layers $V_{\text{int}}$ (with the bottom BLG grounded), we can control the carrier density and displacement field in each BLG, and selectively tune them into different flavor polarization states. More specifically, the carrier densities are given by $n_{\text{top}}=\frac{1}{e}\left(C_{\text{t}}(V_{\text{TG}}-V_{\text{int}})-C_{\text{int}}V_{\text{int}}\right)$ and $n_{\text{bot}}=\frac{1}{e}(C_{\text{b}}V_{\text{BG}}+C_{\text{int}}V_{\text{int}})$, and the displacement fields $D_{\text{top}}/\epsilon_{0}=\frac{1}{2}(C_{\text{t}}(V_{\text{TG}}-V_{\text{int}})+C_{\text{int}}V_{\text{int}})$ and $D_{\text{bot}}/\epsilon_{0}=\frac{1}{2}(-C_{\text{b}}V_{\text{BG}}+C_{\text{int}}V_{\text{int}})$. Here, $e$ is the electron charge, $\epsilon_{0}$ is the vacuum permittivity, and $C_{\text{t}}$, $C_{\text{int}}$ and $C_{\text{b}}$ are the geometric capacitances between the top graphite and top BLG, between the top and bottom BLG layers, and between the bottom BLG and bottom graphite, respectively (see Supplementary Information (SI) S1). Similar to our previous experiment [20], we measure the longitudinal/Hall resistance of both the drive/drag layers $R_{\text{drive/drag}}^{\text{xx/xy}}$. A nontrivial drag signal often provides an unambiguous indication of an interlayer state. Unless otherwise specified, most of the data presented in the main text were acquired with the bottom BLG as the drive layer and the top BLG as the drag layer at temperature $T=250$ mK and magnetic field $B=16$ T.
@@
We first examine the QH phase diagram of the bottom BLG by fixing $V_{\text{TG}}$ to zero and sweeping only $V_{\text{int}}$ and $V_{\text{BG}}$. In this configuration, we effectively use the top BLG as a top gate for the bottom BLG. Fig. 1d shows $R_{\text{xx}}^{\text{drive}}$ as a function of $D_{\text{bot}}$, and the LL filling factor, $\nu_{\text{bot}}=n_{\text{bot}}/n_{0}$, where $n_{0}=eB/h$ is the LL degeneracy per unit area. The range of $\nu_{\text{bot}}$ spans from -2 to 4, covering a substantial portion of the zero-energy LL manifold. Generally, we observe vertical features that are consistent with previous studies. The large regions of zero $R_{\text{xx}}^{\text{drive}}$, shown in deep blue near integer $\nu_{\text{bot}}$, correspond to the integer QH states. The resistive regions in between indicate partially filled LLs. The data show two distinct types of partially filled LLs: one characterized by sharp resistance minima, which are indicative of FQH states with a denominator of 3 and appear mainly between $0<\nu_{\text{bot}}<1$ and $2<\nu_{\text{bot}}<3$. This is characteristic of the $N=0$ LL orbitals in BLG. The other type corresponds to the partially filled $N=1$ LL, which lacks such fractional features with vanishing $R_{\text{xx}}^{\text{drive}}$, and predominantly appears in the range $1<\nu_{\text{bot}}<2$ and $3<\nu_{\text{bot}}<4$.
@@
In addition to the QH states, we also observe several resistive horizontal lines indicative of phase transitions at constant $D_{\text{bot}}$. Prominent examples lie at $D_{\text{bot}}/\epsilon_{0}\approx 0$ across the full $\nu_{\text{bot}}$ range and at $D_{\text{bot}}/\epsilon_{0}\approx\pm 100~\text{mV/nm}$ near $\nu_{\text{bot}}=1$. There are subtler transition lines that become more visible in the numerical derivative of the data (see SI S3). These features originate from ground state rearrangement induced by changes in the QH state hierarchy. In Fig. 1e, we illustrate these features as dashed horizontal lines, together with solid lines that mark the FQH states and the boundaries of integer QH states. This phase diagram matches the previously studied BLG flavor transition maps obtained at comparable magnetic fields [29, 8]. Based on this comparison, we can precisely assign the flavor polarization to each LL segment in Fig. 1e, with the same color coding as Fig. 1c for different valley and orbital flavors. We also note that the phase boundaries between different LLs exhibit zigzag shapes. This irregularity occurs because as $V_{\text{int}}$ changes, the top BLG intermittently enters its own gapped QH states and momentarily stops to act as a gate.
@@
Setting $V_{\text{int}}=0$, we now tune $\nu_{\text{top}}$ and $\nu_{\text{bot}}$ independently via $V_{\text{TG}}$ and $V_{\text{BG}}$ and study the collective phase diagram of the two BLG layers. Fig. 2a and 2b show $R_{\text{xx}}^{\text{drive}}$ and the corresponding $R_{xx}^{\text{drag}}$ over $4\leq\nu_{\text{top}}\leq 4$ and $0\leq\nu_{\text{bot}}\leq 4$. The vertical stripe features are primarily QH states of the drive layer — the bottom BLG. Based on the aforementioned denominator-3 FQH features, we can already infer that the bottom BLG occupies $N=0$ LL for $0\leq\nu_{\text{bot}}\leq 1$ and $2\leq\nu_{\text{bot}}\leq 3$, and $N=1$ LL for $1\leq\nu_{\text{bot}}\leq 2$ and $3\leq\nu_{\text{bot}}\leq 4$. This assignment is confirmed by computing the corresponding $D_{\text{bot}}$ (see SI Fig. S5c) and comparing with the single-layer BLG phase diagram. Similar analysis shows that the top BLG resides in $N=0$ LL for $\nu_{\text{bot}}$ intervals [4,3][-4,-3], [2,1][-2,-1], $[0,1]$, $[2,3]$, and in $N=1$ LL elsewhere. The zigzag shape of the vertical stripes again reflects the quantitative chemical potential change in the top BLG (see discussions in SI S3).
@@
Beyond the intralayer QH features, we also observe interlayer features, highlighted by the eight dashed squares in Fig. 2a and 2b. These features arise only when both BLG layers occupy $N=0$ LL, and they manifest as diamond-shaped features with a diagonal stripe running along lines of constant integer total filling $\nu_{\text{tot}}=\nu_{\text{top}}+\nu_{\text{bot}}$. Along each diagonal line, the number of quasi-holes in one layer equals to the number of quasi-electrons in the other. Furthermore, when $\nu_{\text{top}}>0$ and $\nu_{\text{bot}}>0$, the diagonal stripes appear as resistance minima approaching zero in both $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$, while their $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$ lock to the quantized value $\frac{1}{\nu_{\text{tot}}}\frac{h}{e^{2}}$. These signatures are consistent with the ECs in the previous Coulomb-drag studies of double monolayer graphene [20, 21] and double BLG [18, 19]. Fig. 2c and 2d zoom into the region around $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,2.5)$ at $B=$ 16 T and 25 T, respectively, and Fig. 2e around $(0.5,0.5)$, all clearly revealing a wide region of zero $R_{\text{xx}}^{\text{drag}}$ along the diagonal. An off-diagonal line cut taken along $\nu_{\text{bot}}=\nu_{\text{top}}-2$ (Fig. 2f) shows that exacly at $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,2.5)$, where the EC occurs, $R_{\text{xx}}^{\text{drag}}=R_{\text{xx}}^{\text{drive}}=0$ and $R_{\text{xy}}^{\text{drag}}=R_{\text{xy}}^{\text{drive}}=\frac{1}{3}\frac{h}{e^{2}}$.
@@
What’s also prominent in these diamond-shaped regions are the correlated fractional states of both intra- and interlayer nature. For example, the vertical streaks of $R_{\text{xx}}^{\text{drag}}=0$ at $\nu_{\text{bot}}=\frac{2}{3}$ in Fig. 2c and 2e, as well as the horizontal streaks of $R_{\text{xx}}^{\text{drag}}=0$ at $\nu_{\text{top}}=2+\frac{2}{3}$ in Fig. 2c and $\nu_{\text{top}}=\frac{2}{3}$ in Fig. 2e correspond to intra-layer FQH states that developed independently in each BLG. There are also $R_{\text{xx}}^{\text{drag}}$ peaks that follow fractional-slope lines corresponding to interlayer FQH states, including $\nu_{\text{bot}}+\frac{2(\nu_{\text{top}}-2)}{3}=1$ and $\nu_{\text{bot}}+\frac{3(\nu_{\text{top}}-2)}{5}=1$ in Fig. 2c, and $\nu_{\text{bot}}+\frac{2\nu_{\text{top}}}{3}=1$, $\nu_{\text{bot}}+\frac{\nu_{\text{top}}}{3}=1$ and $\frac{\nu_{\text{bot}}}{3}+\nu_{\text{top}}=1$ in Fig. 2e (see details in SI Fig. S8-S10). At higher magnetic fields, additional intra- and interlayer FQH states emerge. For example, compared to 16 T (Fig. 2c), the data at 25 T (Fig. 2d) presents additional intralayer FQH states at $\nu_{\text{top/bot}}=\frac{1}{3},\frac{3}{5}$, and interlayer FQH states along $\frac{3\nu_{\text{bot}}}{2}+(\nu_{\text{top}}-2)=1$ and $\nu_{\text{bot}}+\frac{3(\nu_{\text{top}}-2)}{2}=1$. These interlayer FQH states with fractional slopes, termed semi-quantized states, have been previously seen in double monolayer graphene [20] and are attributed to electrons binding interlayer vortices via interlayer Coulomb interactions. The fact that these interlayer FQH states appear in $N=0$ LL but not $N=1$ in double BLG mirrors the behavior of double monolayer graphene  [20, 21], suggesting that the interlayer FQH state structure dependends more on the LL wavefunction than on the host material itself.
@@
Despite the rich structure at $V_{\text{int}}=0$, no interlayer states involving N = 1 LLs are seen. This changes when we introduce a finite interlayer bias. Fig. 3a and 3b present $R_{\text{xx}}^{\text{drive}}$ and the corresponding $R_{\text{xx}}^{\text{drag}}$ at $V_{\text{int}}=0.07~\mathrm{V}$. Now the $N=0$ diamond features, again marked by dashed squares, survive only near four locations: $(\nu_{\text{bot}},\nu_{\text{top}})=$ $(0.5,-1.5\pm 2)$ and $(2.5,-1.5\pm 2))$, and disappear near $(0.5,0.5\pm 2))$ and $(2.5,0.5\pm 2))$. Strikingly, new interlayer states emerge when the bottom BLG lies in $1<\nu_{\text{bot}}<2$ or $3<\nu_{\text{bot}}<4$ and the top BLG in $3<\nu_{\text{top}}<-2$ or $1<\nu_{\text{top}}<0$, marked by dashed circles. Each appears as a diagonal minimum in $R_{\text{xx}}^{\text{drive}}$ along an integer $\nu_{\text{tot}}$, accompanied by a non-zero drag signal, confirming their interlayer nature. Based on the QH phase diagram of single BLG, we can identify their locations as both layers occupying $N=1$ LL. The two zoomed-in views overlaying Fig. 3a highlight the visual distinction between the previous $N=0$ and the new $N=1$ interlayer states. The $N=0$ state on the left has a characteristic wide diagonal zero $R_{\text{xx}}^{\text{drive}}$ feature that cuts into the resistive region from one side and is surrounded by other fractional states. By contrast, the $N=1$ state on the right manifests itself as a narrow zero $R_{\text{xx}}^{\text{drive}}$ line that cuts through the resistive region from both sides, without other prominent features in its surroundings. Raising $V_{\text{int}}$ to $0.1~\mathrm{V}$ reveals still more $N=1$ interlayer states (see SI Fig. S7).
@@
Although visually distinct, the $N=1$ interlayer states show transport hallmarks of ECs: $R_{\text{xx}}^{\text{drag/drive}}=0$ and $R_{\text{xy}}^{\text{drag/drive}}=\frac{1}{\nu_{\text{tot}}}\frac{h}{e^{2}}$. Fig. 3c and 3e display $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drag}}$ for the $N=1$ interlayer state near $(\nu_{\text{bot}},\nu_{\text{top}})$ = $(1.5,0.5)$ at $V_{\text{int}}=0.1~\mathrm{V}$. This interlayer state exhibits as a thin line of $R_{\text{xx}}^{\text{drag}}$ dip within a narrow resistive strip that runs along the diagonal, and as a peak in $R_{\text{xy}}^{\text{drag}}$ along the same line. Examining the off-diagonal linecut along $\nu_{\text{bot}}=\nu_{\text{top}}-1$, we observe that at $(\nu_{\text{bot}},\nu_{\text{top}})$ = $(1.5,0.5)$, $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$ approach the quantization value $\frac{1}{2}\frac{h}{e^{2}}$ (Fig. 3d), while $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$ drop to 0 (Fig. 3f), consistent with an interlayer EC. Note that this $N=1$ interlayer state appears for $V_{\text{int}}=0.1$  V, but is absent in $V_{\text{int}}=0.07$ V (Fig. 3b).

### v2

@@ Abstract
Electrons moving in two-dimensional (2D) space and under a strong magnetic field fall into discrete Landau levels (LLs), indexed by an orbital quantum number $N=0,1,2$..., with the corresponding spatial wavefunction becoming increasingly nodal as $N$ increases. The ultraflat electronic dispersion creates an ideal environment for Coulomb interactions to dominate and induce various many-body ground states, such as the Wigner crystal [8], charge density wave [9], and most notably, the fractional quantum Hall (FQH) effect [34]. The orbital character is critical to interaction physics because the Haldane pseudopotential for LL wavefunctions varies with $N$ [11]. In particular, even-denominator FQH states at half-filling have been predominantly observed in the $N=1$ LL [35, 29, 5, 40, 20, 32], and are believed to be described by a class of non-Abelian states including the Moore-Read Pfaffian state [27]. Experimentally, access to different LLs effectively serves as a powerful tuning knob for electron-electron interactions, fostering exotic many-body phases.
@@
Quantum Hall (QH) physics can be further enriched by placing two electrically isolated, parallel layers of 2D electron systems close to each other so that electrons in one layer interact solely through Coulomb interactions with those in the other. This double-layer structure enables more complex multi-component states by introducing a new pseudospin—the layer index, and provides unique experimental access to coherent QH states by allowing for probing the two spatially separated pseudospin components, something impossible for a spin or valley based multi-component state. In particular, the multi-component Halperin (111) state  [12] composed of layer pseudospins maps to an interlayer exciton condensate (EC) formed from electrons and holes in the partially filled LLs in the opposite layers [15, 7]. Such interlayer ECs were first observed in GaAs double quantum wells [16, 28], and later realized in double layer structures of monolayer or Bernal-stacked bilayer graphene (BLG) [25, 21, 23, 19]. The exceptional flexibility and tunability of these double-layer graphene structures further enabled observation of the BEC-BCS crossover [24], novel interlayer FQH states [23, 19] and exciton solids [37].
@@
However, these interlayer correlated states have been observed only when both layers are in the $N=0$ LL; clear $N=1N=1$ interlayer states remain elusive. Since $N=1$ can host a different hierarchy of single-layer FQH states, it could likewise support qualitatively new interlayer states in double-layer systems. Furthermore, once an EC between two $N=1$ LLs is realized, the transition between this interlayer coherent EC and individual-layer even-denominator FQH states could be a highly nontrivial topological phase transition [31, 39]. Previously, there have been reports of gapped states in 2H-stacked bilayer $\mathrm{WSe_{2}}$ [33] or twisted double bilayer graphene [22] that are consistent with interlayer ECs in higher LLs. However, while the absence of a spacer helps achieve strong interlayer coupling, it prevents probing each layer pseudospin separately, making it impossible to conclusively identify the interlayer coherence of EC states. In this Letter, we present the first observation of quantized Coulomb-drag signals between the $N=1$ LLs in a double bilayer graphene device, providing key signatures of EC formation in higher LLs.
@@
Fig. 1a shows the device image, and Fig. 1b illustrates its layer configuration. Two BLG layers are separated by a 2.5 nm thick hBN spacer and have separate gold contacts. The top and bottom graphite gates control the channel region, while the silicon back gate and additional contact gates ensure good equilibration between the channel region and gold contacts.
@@
The zero-energy LL of BLG consists of eight quasi-degenerate flavors: spin up and down, valley $K$ and $K^{\prime}$, and an orbital quantum number, $N=0$ and 1 [26]. Fig. 1c highlights the valley and orbital flavors. The $N=0$ orbital wavefunction is similar to the lowest LL in a conventional quadratic-band system like GaAs. By contrast, the $N=1$ orbital in BLG is an admixture of the conventional lowest and second LL wavefunctions [26]. Depending on the valley flavor, these wavefunctions are layer-polarized within BLG, illustrated in Fig. 1c. Previous studies have shown that next-order energy anisotropies and Coulomb interactions in BLG induce an energy hierarchy among the eight flavors, leading to their sequential filling as carrier density increases [14, 20, 40]. A displacement field further modifies the hierarchy by shifting the relative potential within BLG and altering the energetics according to the varying degree of layer polarization.
@@
In our experiment, by applying DC voltages to the top and bottom graphite gates, $V_{\text{TG}}$ and $V_{\text{BG}}$, as well as between the two BLG layers $V_{\text{int}}$ (with the bottom BLG grounded), we can control the carrier density $n_{\text{top/bot}}$ and displacement field $D_{\text{top/bot}}$ in each BLG (see Supplemental Material, hereafter SM [1], Sec. S1), and selectively tune them into different flavor polarizations. Similar to our previous experiment [23], we measure longitudinal/Hall resistances in both drive and drag layers $R_{\text{drive/drag}}^{\text{xx/xy}}$. A nontrivial drag signal often provides a direct indication of an interlayer state. Unless otherwise specified, most of the data presented in the main text were taken with the bottom BLG as the drive layer and the top BLG as the drag layer at temperature $T=250$ mK and magnetic field $B=16$ T.
@@
We first examine the QH phase diagram of the bottom BLG by fixing $V_{\text{TG}}$ to zero and sweeping only $V_{\text{int}}$ and $V_{\text{BG}}$, so that the top BLG effectively serves as a top gate for the bottom BLG. Fig. 1d shows $R_{\text{xx}}^{\text{drive}}$ as a function of $D_{\text{bot}}$, and the LL filling factor, $\nu_{\text{bot}}=n_{\text{bot}}/n_{0}$, where $n_{0}=eB/h$ is the LL degeneracy per unit area. The range of $\nu_{\text{bot}}$ spans from -2 to 4, covering a substantial portion of the zero-energy LL manifold. Generally, we observe vertical features that are consistent with previous studies. The large regions of zero $R_{\text{xx}}^{\text{drive}}$ near integer $\nu_{\text{bot}}$, shown in deep blue, correspond to integer QH states, while the intervening resistive regions indicate partially filled LLs. Two types of partially filled LLs are evident. One has sharp resistance minima indicative of FQH states with a denominator of 3, appearing mainly for $0<\nu_{\text{bot}}<1$ and $2<\nu_{\text{bot}}<3$. This is characteristic of the $N=0$ LL orbitals in BLG. The other type lacks fractional features and predominantly appears in $1<\nu_{\text{bot}}<2$ and $3<\nu_{\text{bot}}<4$, consistent with occupation of the $N=1$ LL orbital.
@@
In addition, there are several resistive horizontal lines consistent with displacement field-induced flavor hierarchy transitions. Prominent examples lie near $D_{\text{bot}}/\epsilon_{0}\approx 0$ across the full $\nu_{\text{bot}}$ range and near $D_{\text{bot}}/\epsilon_{0}\approx\pm 100~\text{mV/nm}$ around $\nu_{\text{bot}}=1$. Some subtler transition lines become clearer in the numerical derivative of the data (SM, Sec. S3). In Fig. 1e, we mark these transitions as dashed horizontal lines, together with solid lines indicating the FQH states and boundaries of integer QH states. This phase diagram matches the established BLG flavor transition maps obtained at comparable magnetic fields [14, 40], allowing us to precisely assign the flavor polarization to each LL segment in Fig. 1e (same color coding as Fig. 1c). We also note that the phase boundaries between different LLs exhibit zigzag shapes. This irregularity occurs because as $V_{\text{int}}$ changes, the top BLG intermittently enters its own gapped QH states and momentarily stops to act as a gate.
@@
Setting $V_{\text{int}}=0$, we now tune $\nu_{\text{top}}$ and $\nu_{\text{bot}}$ independently via $V_{\text{TG}}$ and $V_{\text{BG}}$ and study the collective phase diagram of the two BLG layers. Fig. 2a,b show $R_{\text{xx}}^{\text{drive}}$ and the corresponding $R_{xx}^{\text{drag}}$ over $4\leq\nu_{\text{top}}\leq 4$ and $0\leq\nu_{\text{bot}}\leq 4$. The vertical stripe features are primarily QH states of the drive layer — the bottom BLG. Based on the aforementioned denominator-3 FQH features, we can already infer that the bottom BLG occupies $N=0$ LL for $0\leq\nu_{\text{bot}}\leq 1$ and $2\leq\nu_{\text{bot}}\leq 3$, and $N=1$ LL for $1\leq\nu_{\text{bot}}\leq 2$ and $3\leq\nu_{\text{bot}}\leq 4$. This assignment is confirmed by computing the corresponding $D_{\text{bot}}$ (SM, Fig. S9c) and comparing with the single-layer BLG phase diagram. Similar analysis shows that the top BLG resides in $N=0$ LL for $\nu_{\text{top}}$ intervals [4,3][-4,-3], [2,1][-2,-1], $[0,1]$, $[2,3]$, and in $N=1$ LL elsewhere. The zigzag shape of the vertical stripes again reflects the quantitative chemical potential change in the top BLG (see discussions in SM, Sec. S3).
@@
Beyond these intralayer QH features, we also observe interlayer features, highlighted by the eight dashed squares. They arise only when both BLG layers occupy $N=0$ LL, and appear as diamond-shaped features with a diagonal stripe running along lines of constant integer total filling $\nu_{\text{tot}}=\nu_{\text{top}}+\nu_{\text{bot}}$, where the number of quasi-holes in one layer equals the number of quasi-electrons in the other. For $\nu_{\text{top}}>0$ and $\nu_{\text{bot}}>0$, the diagonal stripes appear as resistance minima approaching zero in both $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$, while their $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$ lock to the quantized value $\frac{1}{\nu_{\text{tot}}}\frac{h}{e^{2}}$. These signatures are consistent with the ECs in the previous Coulomb-drag studies of double monolayer graphene [23, 19] and double BLG [25, 21]. Fig. 2c,d zoom into the region around $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,2.5)$ at $B=$ 16 T and 25 T, respectively, and Fig. 2e around $(0.5,0.5)$, all clearly revealing a wide region of zero $R_{\text{xx}}^{\text{drag}}$ along the diagonal. An off-diagonal line cut taken along $\nu_{\text{bot}}=\nu_{\text{top}}-2$ (Fig. 2f) shows that exactly at $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,2.5)$, where the EC occurs, $R_{\text{xx}}^{\text{drag}}=R_{\text{xx}}^{\text{drive}}=0$ and $R_{\text{xy}}^{\text{drag}}=R_{\text{xy}}^{\text{drive}}=\frac{1}{3}\frac{h}{e^{2}}$. Note that the filling factors in the large-area maps (Fig. 2a,b) are estimated ignoring chemical-potential effects, and thus indicate only approximate state locations. In the zoomed-in maps and line cuts (Fig. 2c–e), where accurate $\nu$ is required, filling factors are calibrated against known fractional QH features (see SM, Sec. S1 for details).
@@
These diamond-shaped regions also host correlated fractional states of both intra- and interlayer nature. The vertical and horizontal streaks near fractional $\nu_{\text{top/bot}}$ in Fig. 2c,e correspond to intra-layer FQH states developing independently in each BLG. In addition, there are straight lines of resistance peak/dip features following fractional slopes. They are interlayer FQH states termed semi-quantized states, previously observed in double monolayer graphene [23] and attributed to electrons binding interlayer vortices via interlayer Coulomb interactions. We defer the detailed analysis of these states to SM, Sec. S5. The fact that these interlayer FQH states appear in $N=0$ LL but not in $N=1$ in double BLG mirrors the behavior of double monolayer graphene  [23, 19], suggesting that the interlayer FQH state structure depends more on the LL wavefunction than on the host material itself.
@@
Despite the rich structure at $V_{\text{int}}=0$, we observe no interlayer states involving N = 1 LLs. This changes when we introduce a finite interlayer bias. Fig. 3a,b show $R_{\text{xx}}^{\text{drive}}$ and $R_{\text{xx}}^{\text{drag}}$ at $V_{\text{int}}=0.07~\mathrm{V}$. Now the $N=0$ diamond features, again marked by dashed squares, survive only near four locations: $(\nu_{\text{bot}},\nu_{\text{top}})=$ $(0.5,-1.5\pm 2)$ and $(2.5,-1.5\pm 2)$. Strikingly, new interlayer states emerge when the bottom BLG lies in $1<\nu_{\text{bot}}<2$ or $3<\nu_{\text{bot}}<4$ and the top BLG in $3<\nu_{\text{top}}<-2$ or $1<\nu_{\text{top}}<0$ (dashed circles). Each appears as a narrow diagonal minimum in $R_{\text{xx}}^{\text{drive}}$ along an integer $\nu_{\text{tot}}$, accompanied by a non-zero drag signal, confirming their interlayer nature. Using the single-layer BLG phase diagram, we can identify their locations as both layers occupying $N=1$ LL. The zoomed-in views overlaying Fig. 3a highlight a clear visual contrast: the $N=0$ state on the left has a characteristic wide diagonal zero $R_{\text{xx}}^{\text{drive}}$ feature that cuts into the resistive region from one side and is surrounded by other fractional states, whereas the $N=1$ state on the right appears as a thin zero $R_{\text{xx}}^{\text{drive}}$ line cutting through the resistive region with few nearby features. Raising $V_{\text{int}}$ to $0.1~\mathrm{V}$ reveals still more $N=1$ interlayer states (SM, Fig. S11).
@@
Although visually distinct, the $N=1$ interlayer states exhibit transport hallmarks of ECs: $R_{\text{xx}}^{\text{drag/drive}}=0$ and $R_{\text{xy}}^{\text{drag/drive}}=\frac{1}{\nu_{\text{tot}}}\frac{h}{e^{2}}$. Fig. 3c,e show $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drag}}$ for the $N=1$ interlayer state near $(\nu_{\text{bot}},\nu_{\text{top}})$ = $(1.5,0.5)$ at $V_{\text{int}}=0.1~\mathrm{V}$. It appears as a thin line of $R_{\text{xx}}^{\text{drag}}$ dip within a narrow resistive strip that runs along the diagonal, and as a peak in $R_{\text{xy}}^{\text{drag}}$ along the same line. A line cut along $\nu_{\text{bot}}=\nu_{\text{top}}-1$ shows that exactly at $(1.5,0.5)$, $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$ approach the quantization value $\frac{1}{2}\frac{h}{e^{2}}$ (Fig. 3f), while $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$ drop to 0 (Fig. 3d), consistent with an interlayer EC. Note that this $N=1$ state is present at $V_{\text{int}}=0.1$ V but absent at $V_{\text{int}}=0.07$ V (Fig. 3b).